OpenStudy (anonymous):

Okay...... So P=1/2(x+y)+2+2w

OpenStudy (radar):

O.K. What is the problem?

OpenStudy (anonymous):

solve 4 p. whoopsies!

OpenStudy (radar):

In terms of x, y and w, it is already solved.

OpenStudy (radar):

Is there more information like some values for x, y, w etc.?

OpenStudy (anonymous):

yes! io totally 4got 2 post tht! Find P, when x = 10, y = 14, z = 18, and w = 20.

OpenStudy (radar):

O.K then it would be: P=12+2+40=54

OpenStudy (radar):

1/2(x+y)= 1/2(10+14)=12 2=2 2w=2(20)=40

OpenStudy (radar):

Put em together you got P=54

OpenStudy (radar):

I did not see a z, so that must be for a another problem.

OpenStudy (anonymous):

oops! sorry! i totally messed up on this! it's P=1/2(x+y)+z+2w